If a firm plans to issue new stock, flotation costs (investment bankers' fees) should not be ignored. There are two approaches to use to account for flotation costs. The first approach is to add the sum of flotation costs for the debt, preferred, and common stock and add them to the initial investment cost. Because the investment cost is increased, the project's expected rate of return is reduced so it may not meet the firm's hurdle rate for acceptance of the project. The second approach involves adjusting the cost of common equity as follows: The difference between the flotation-adjusted cost of equity and the cost of equity calculated without the flotation adjustment represents the flotation cost adjustment. Quantitative Problem: Barton Industries expects next year's annual dividend, D1, to be $1.50 and it expects dividends to grow at a constant rate g = 4.5%. The firm's current common stock price, P0, is $21.00. If it needs to issue new common stock, the firm will encounter a 5.3% flotation cost, F. Assume that the cost of equity calculated without the flotation adjustment is 12% and the cost of old common equity is 11.5%. What is the flotation cost adjustment that must be added to its cost of retained earnings? Do not round intermediate calculations. Round your answer to two decimal places. % What is the cost of new common equity considering the estimate made from the three estimation methodologies? Do not round intermediate calculations. Round your answer to two decimal places.
Answer #1

a) D1=$1.50, g=4.5%, P0 = $21, Flotation Cost F = 5.3%, Cost of equity without flotation cost = 12%, Cost of old common equity = 11.5%

Required return new equity = D Po*(1-F) +9 = 1.5 / (21*(1-0.053) + 0.045 = 0.075426+0.045 = 0.1204=12.04%

Flotation adjustment cost = required return - required return without adjustment

Flotation adjustment cost = 12.04% -12% = 0.04%

b) Cost of new equity with adjustment = Cost of new equity+flotation cost adjustment= 11.5%+0.04% = 11.54%